This word problem is stumping me, can anybody help?
1. Some students planned for a get-together. The budget for the food was $500. Five of the students failed to come therefore the cost of the food for each member increased by $5. How may students attended the get together?

Answer by scott8148(6628)

let x="number attending", so x+5="original number"

"Five of the students failed to come therefore the cost of the food for each member increased by $5"
__ (500/(x+5))+5=500/x

multiplying by x(x+5) __ 500x+5x^2+25x=500x+2500

subtracting 500x and dividing by 5 __ x^2+5x=500 __ subtracting 500 __ x^2+5x-500=0

factoring __ (x+25)(x-20)=0 __ x+25=0 and x-20=0